OpenStudy (anonymous):

Find the point(0,b) on the y-axis that is equidistant from the points (3, 3) and (6, -5). b=

OpenStudy (ribhu):

may i??

OpenStudy (anonymous):

May you help? Is that what your asking?

OpenStudy (ribhu):

yeah!

OpenStudy (anonymous):

sure, be my guest xp

OpenStudy (ribhu):

let the point be (0,b)

OpenStudy (ribhu):

calculate the distance between ( 0,b) and (3,3) sqrt ((3-0)^2+(3-b)^2)

OpenStudy (ribhu):

similarly between (0,b) and (6,-5) which is sqrt ((6-0)^2+(-5-b)^2)

OpenStudy (ribhu):

since the point is equidistant from both the points so the distances must be equal. 9+9+b^2-6b = 36+25+b^2+10b; thiscan be solved for b

OpenStudy (anonymous):

ok thnx, i get it now, ill solve for b now.

OpenStudy (anonymous):

ok i got 0= 43 + 16b -2b^2 i got stuck

OpenStudy (ribhu):

-18 =16b

OpenStudy (anonymous):

how did you get that?

OpenStudy (ribhu):

only solving the equation which i wrote

OpenStudy (anonymous):

never mind i see what i did wrong, i put -b^2 instead of +b^2

OpenStudy (anonymous):

i got -43 = 16b

OpenStudy (ribhu):

sorry u r correct

OpenStudy (anonymous):

aha.....thnx for the help XD